OpenStudy (anonymous):

One number is 5 more than another number. If you add 12 to 5 times the first number, the result is 6 times the second number.

OpenStudy (anonymous):

y=x+5 5y+12=6x Do you know what to do from there?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

Okay lets get y=x+5 . Logically, y equals x plus 5. So plug this in for y into the *other* equation. 5(x+5)+12=6x Distribute. 5x+25+12=6x Simplify 5x+37=6x subtract 37 5x=6x-37 Get the x's on the same side (left) Minus 6 x then. -x=-37 Divide by -1 x=37 Plug 37 in for x . And solve for your Y :]

OpenStudy (anonymous):

y=37+5 (came from y=x+5, with our 37 plugged in for x) y=42